OpenStudy (anonymous):

A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose all answers that are correct. A. −7x + y = −44 B. 7x – y = 44 C. (y − 9) = 7(x + 5) D. (y + 9) = 7(x − 5)

OpenStudy (shinalcantara):

using the point slope form you'll have: (y-y1) = m(x-x1) y - (-9) = 7(x -5) y + 9 = 7(x - 5) or y + 9 = 7x - 35 7x - y = 9 + 35 7x - y = 44

OpenStudy (shinalcantara):

or multiplying the whole equation with -1 7x - y = 44 -7x + y = -44